OpenStudy (anonymous):

3x-2/2x^2-5x-3 (-) 1/x-3 subtract

OpenStudy (anonymous):

\[\frac{3x-2}{2x^2-5x-3}-\frac{1}{x-3}\]?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

alright so first you want to factor the denominator of the first part, do you know how to do this?

OpenStudy (anonymous):

No i don

OpenStudy (anonymous):

alright so first you want to multiply your leading coefficient by -3 2*-3=-6 now you need to figure out all the factors of -6 1*-6 2*-3 3*-2 6*-1 now you need to figure out which add up to your middle coefficient hwich is -5 -6+1=-5 so your numbers are -6 and 1 replace the middle term with these two terms like this \[2x^2-6x+1x-3\] now you can group both of these \[(2x^2-6x)+(x-3)\] pul out 2x from th first equation you get \[2x(x-3)+1(x-3)\] your answer is going to be the function in parenthesis and the one outside soo (2x+1)(x-3)

OpenStudy (anonymous):

so now you have \[\frac{3x-2}{(2x+1)(x-3)}-\frac{1}{x-3}\]

OpenStudy (anonymous):

now when subtracting to fractions you must always create a common denominator...looking at this, your second denominator is missing (2x+1) so multiply it by 1 (2x+1/2x+1) \[\frac{3x-2}{(2x+1)(x-3)}-\frac{1}{x-3}*\frac{2x+1}{2x+1}\] multiplying it out you get \[\frac{3x-2}{(2x+1)(x-3)}-\frac{2x+1}{(x-3)(2x+1)}\] you can now combine like the fractions to get \[\frac{3x-2-2x-1}{(2x+1)(x-3)}\] add like terms you get \[\frac{x-3}{(2x+1)(x-3)}\] you can cancel out the x-3's so you're left with \[\frac{1}{2x+1}\]